Which of the following is true regarding the national income accounts?
a. GDP>GNP
b. NNP>GNP
c. Personal Income>National income
d. Personal Income>NNP
e. Personal Income>Disposable Personal Income
1
Expert's answer
2020-09-21T08:20:08-0400

"ANSWER"

Only e is true.


"SOLUTION"

a) "(GDP > GNP):" FALSE

GDP maybe greater than or less than GNP. The relationship depends on the value of net property factor income from abroad.


GNP = GDP + Net property factor income from abroad

= GDP + NPIFA


GNP adds income from abroad to GDP. This income is in the form of profits, dividends, interests, and other incomes received by the nationals of the country from abroad in excess of the same income received by nationals of other countries in our domestic economy. Now, in the event that NPIFA is negative, GDP > GNP.


NPIFA = Income from abroad - Income paid abroad. Therefore, if income paid abroad exceeds income received by nationals from abroad, NPIFA becomes negative and GDP will exceeds GNP otherwise GNP exceeds GDP.


b) "(NNP > GNP):" FALSE

GNP exceeds NNP.


Net national product (NNP) is Gross national product (GNP) less Depreciation.

NNP = GNP - depreciation

 Therefore, NNP is less than GNP because it is GNP after depreciation has been deducted. Thus, the correct statement is GNP > NNP.


c) "(PI > NI):" FALSE

Personal income maybe greater than or less than than national income.


Personal Income (PI) is national income Plus net income received but not earned such as interest on government debt, and transfer payments such social security, Minus net income earned but not received, for example, retained corporate earnings.

PI = Nl + Income received but not earned - Income earned but not received. Thus, the relationship between PI and NI is dependent upon the sizes of incomes received but not earned and income earned but not received.

 

d) "(PI > NNP):" FALSE

PI may be greater than or less than NNP.


PI = NI + net income received but not earned - net income earned but not received. Where, NI = NNP - depreciation.

=> PI = (NNP - depreciation) + net income received but not earned - net income earned but not received.

=> PI = NNP + net income received but not earned - (net income earned but not received + depreciation).


Thus, the sizes of net income earned but not received, depreciation, and net income received but not earned control the relationship between PI and NNP. Hence, PI can be greater than or less than NNP.


e) "(PI > PDI):" TRUE

PI = PDI + personal taxes, hence true.


Disposable personal income = Personal Income - Personal Taxes. Thus, PI exceeds PDI by the amount of personal taxes.